Operations Report – April 18th

Crew 280 Operations Report 18-04-2023
SOL: 2
Name of person filing report: Neus Sabaté – Mission Specialist (Assisted by Cesca Cufí – Crew Engineering)
Non-nominal systems: robotic observatory
Notes on non-nominal systems: Replacement parts are scheduled to be installed during crew 261 rotation.
Spirit rover used: yes
Hours: 216.7
Beginning charge: 100%
Ending charge: 88%
Currently charging: no
Opportunity rover used: no
Hours: 120.7
Beginning charge: –
Ending charge: –
Currently charging: no
Curiosity rover used: yes
Hours: 229.8
Beginning charge: 100%
Ending charge: 92%
Currently charging: no
Perseverance rover used: no
Hours: 261.1
Beginning charge: –
Ending charge: –
Currently charging: no
General notes on rovers: non
Summary of Hab operations:
Water use: can not be estimated because water tank has been partially refilled
Water (static tank): 280 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: yes
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: spacesuits 1,2,5 and 6 have been used during the EVA. All nominal.
Radios: all operative during EVA
All radio and spacesuit batteries checked before EVA: ok.
Radio batteries charged after EVA.
Spacesuit batteries plugged.
Summary of GreenHab operations:
Water use: 14 gallons
Heater: off
Supplemental light: off
Harvest: 191 gr of cucumber
Radios were found discharged: the jack plug of charger was half disconnected so the radios were not properly charged. This has already been solved.
Summary of Science Dome operations:
Laia Ribas (Green Hab Officer) prepared the setups for her experiments.
Neus Sabaté (Mission Specialist) brought rock samples from EVA and prepared the setup for further analysis of electroactive iron contents.
Dual split: off
Summary of RAM operations:
Cesca Cufí (Crew Engineer) used the RAM to start mounting her experiment setup (homemade biological safety cabinet).
Summary of any observatory issues: none
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues:
Gas stove knobs dammaged: the gas stove knobs are damaged and they don’t fit properly the valves axis, resulting in a serious hazard. Because of this reason, there was a gas leak during cooking lunch, while during the EVA. The two cooks detect immediately the leak but could not stop it immediately as the knob was not working and the other knobs neither. The leakage was stopped several minutes after by the crew engineer (cooking at this time) that turned off the knob axis using a pliers that she had in the Hab, while one crew member went to the RAM to pick a gas sensor to monitor the concentration of gas. After the EVA, to ensure the safety, the Crew Engineer explained the issue to all crew members. The HSO has recommended limiting the use of the stove to the single burner that has an actual functional knob while we find a way to repair the rest. A gas detector has been placed in the kitchen to measure the gas concentration in case we need to use several burners, to ensure all burners are off when finishing cooking. All crew members are aware and know the procedure to follow.
Questions, concerns and requests to Mission Support:
The gas detector that has been used to detect the gas leak from the kitchen stove has a very low battery level. Also the second one we found in RAM. We have looked for replacement but could not find AAA batteries in the RAM.
In case Mission Control refuses the crew repairing/substituting the knobs, we require spare parts to replace all of them (x4). We can provide the knob measures and pictures.

Copyright © The Mars Society. All rights reserved. | Main Site